# Anatomy of an ionized bubble: NIRCam grism spectroscopy of the double-peaked Lyman- emitter COLA1 and its environment

### v1

@@ 5 The environment of COLA1: a normal over-density for its UV and [O III] luminosity
During the EoR, galaxy over-densities are associated to an increased Ly transmission (e.g., Endsley & Stark, 2022; Leonova et al., 2022; Kashino et al., 2023; Witstok et al., 2023), as these over-densities are more likely to form ionized bubbles, according to simulations (Ocvirk et al., 2020; Hutter et al., 2021; Qin et al., 2022; Lu et al., 2023). As sight-lines that allow for the detection of blue Ly photons at are expected to be extremely rare (e.g., Gronke et al., 2021), one could therefore expect that galaxies that show a double-peaked Ly line are surrounded by an unusually high over-density.
To address exactly this question, we analyse the over-density around COLA1 and compare it with results from mock observations of a simulation, with the methodology we described in Sect. 3.4. We identify five [O III] emitters in our sample that are close to COLA1 in redshift space (; three of which have erg s-1). These galaxies are all considerably fainter than COLA1, spanning UV luminosities M to M. Notably, as shown in Fig. 11 the sky position of all those objects fall within COLA1’s inferred ionized bubble radius (Matthee et al., 2018; Mason & Gronke, 2020), assuming a spherically symmetrical bubble, meaning that they could potentially be located within the spatial extent of the bubble. By comparing the number of identified sources with the expectation from pure random mocks, we find that COLA1 is located in a mild over-density (). The typical halo mass for galaxies in our mocks that are surrounded by a similar number of galaxies as COLA1 is .
@@
This analysis shows that, while we do identify an over-density around COLA1, its amplitude is typical for any galaxy that is similarly UV bright at . Thus, there is no strong indication that the detection of the double-peaked Ly emission is due to an exceptionally large over-density of galaxies. Also, we remark that Fig. 12 shows that the possible number of neighbours that one could detected around a similarly UV bright galaxy with a similar emission-line survey on a field of arcmin2 has a wide distribution. This displays the wide variety of environments that could host similarly UV bright (; see e.g., Bouwens et al., 2015) galaxies, and therefore stressing the importance of taking cosmic variance of the large scale structure into account when interpreting Ly observations of galaxies in the context of reionization.
@@ 6 Discussion
In this section, we discuss the results presented in Sections 4 and 5. We comment the measured properties of COLA1 in comparison with other SFGs and LyC leakers in the literature. Our results confirm the systemic redshift of COLA1, which means it must reside inside an ionized bubble. We discuss the role of COLA1 and the galaxies in its proximity in ionizing their environment.
@@ 6.1 COLA1 as an ionizing photon leaker
COLA1 presents an extremely blue UV slope (see Table 2), far bluer than the typical values of the C1F+EIGER sample and the best fit in Bouwens et al. (2014) for Lyman-break galaxies at (see Fig. 8). Topping et al. (2023) presented a sample of 44 galaxies with steep UV slopes (), and found that the SEDs of such galaxies could be explained with density-bounded models with high escape fractions () (see also Topping et al., 2022). Further, Kim et al. (2023) found a relation between very steep and LyC leaking regions of a lensed galaxy. Indeed, our estimation based on Chisholm et al. (2022) using suggests a very high (see Table 3).
The rest-frame [O III]+H EW of COLA1 is higher than the typical values for sources with similar UV magnitudes, showing COLA1 is experiencing an intense burst of star formation. The intrinsic [O III] and H luminosities are remarkably high, and . The high [O III] luminosity can also be related to an elevated star formation (Villa-Vélez et al., 2021), as well as the high ratio of (Dickey et al., 2016).
COLA1 presents significantly higher than the average C1F+EIGER values (see Fig. 9), consistent with the measurements for high- SFGs in the literature (e.g., Bouwens et al., 2016; Nakajima et al., 2018; Saxena et al., 2023a; Simmonds et al., 2023, 2024) and also known low- LyC leakers (e.g., Izotov et al., 2016). Izotov et al. (2021) showed that low- compact SFGs are likely analogs of high- SFGs. More recently, compact SFGs have been found at with similar sizes to COLA1, presenting signs of LyC leakage (Mascia et al., 2023b, a).
As stated in Sect. 4.5, the direct measurement of H allows for an estimation of the Ly escape fraction, yielding a significantly high value of (Ly) . This escape fraction is high in comparison with the typical values found in high redshift surveys (e.g., (Ly) %, , Roy et al. 2023; (Ly) %, , Tang et al. 2023), although other objects with similar Ly escape fractions have recently been found (e.g., Chen et al., 2023; Saxena et al., 2023b).
Almost every observed LyC leaker in the literature presents higher than the average at similar redshifts (e.g, Naidu et al., 2020). This is also the case of COLA1. Comparably, the galaxies Ion2 and Ion3 show fairly higher than the average at their respective redshifts, also presenting complex, multi-peaked Ly profiles (Vanzella et al., 2016, 2018).
@@ 6.2 Size of the ionized bubble
The size of the implied ionized region in the line of sight to COLA1 can be derived by the cut-off in the blue wing of the Ly line at km s-1 (see Fig. 4). This cut-off is particularly abrupt compared to the red peak, and defines the the line-of-sight extent in which the IGM is optically thin to Ly photons (so-called proximity zone; Mason & Gronke, 2020). The blue-peak cut-off seen in the Ly profile of COLA1 yields a proximity zone of pMpc (Matthee et al., 2018). Mason & Gronke (2020) argue that the total size of the ionized region extends beyond the proximity zone as there may be significant damping wing absorption. They find that a bubble with a radius of at least pMpc is needed in order to explain the observed velocity offset of COLA1’s blue Ly peak, assuming a residual neutral fraction of and defining the minimum observable velocity offset where the IGM transmission of Ly photons drops to 10%.
As discussed in Mason & Gronke (2020), for a central source to be able to ionize a bubble with the required size pMpc, it needs a steep UV slope (), large escape fraction of ionizing photons (%) and an under-dense line-of-sight gas. Our results show that COLA1 aligns with the first two conditions, but we do not find a strong indication for a specifically under-dense line-of-sight (see Sect. 5). However, the estimated UV slope of COLA1 is steeper than the values considered in Mason & Gronke (2020), and its extreme value could explain this discrepancy as it implies a higher ionizing photon production efficiency. A more detailed analysis of the ionization conditions of the bubble is needed in order to address this topic, for example using Ly line measurements from multiple neighbouring galaxies and by detailed comparisons to radiative hydrodynamical simulations (e.g. Gronke et al., 2021), which is beyond the scope of this work. In Sect. 6.3 below we discuss the role of COLA1 in ionizing such bubble.
@@ 6.3 COLA1 ionizing its surroundings
We investigate the contribution from COLA1 and the surrounding galaxies in the C1F group to the ionization of a pMpc bubble needed to explain COLA1’s Ly profile (see Sect. 6.2). The capability of galaxies of ionizing their medium directly depends on their LyC escape fraction. We estimate (LyC) for our sample with F200W detection (4 out of those 5 objects are detected in this band) using the relation with (UV) in Naidu et al. (2020), as described in Sect. 4.5, for which we get (LyC) and %. For these estimates we have assumed the average dust correction introduced in Sect. 6.1.
@@
Equation 6 in Mason & Gronke (2020) predicts the radius of an ionized bubble powered by a central source as a function of an age parameter for a simple spherical geometry around the ionizing source. Using this equation, and following Witstok et al. (2023), we compute the radius of an ionized bubble powered by COLA1 for different values of the LyC escape fraction (Fig. 13). We have assumed an average hydrogen density cm-3 (Mason & Gronke, 2020) and an ionizing photon production of s-1 (Schaerer et al., 2016). For accounting the contribution from the other galaxies, in Fig. 13 we have conservatively assumed a high fiducial value of (LyC) 888For the computation of for the galaxies other than COLA1 we adopted ( Hz erg-1; Saxena et al., 2023a), since a detection of H is not available for some of the objects.. According to this estimate, COLA1 alone could ionize a 0.7 pMpc bubble in a reasonable span of time ( Myr; see Witstok et al., 2023; Whitler et al., 2023) for escape fractions as low as , assuming it maintained the observed star-formation activity. The other four detected galaxies at make only a minor contribution in ionizing the bubble when combined with COLA1, specially for high COLA1 escape fractions (dotted lines in Fig. 13). For instance, for COLA1 LyC escape fraction of % (30%, 20%), the time required to ionize the bubble is Myr (58, Myr), with COLA1 contributing to (87%, ) of the total ionizing photon budget of the system detected in C1F.
@@
Nevertheless, systems with extremely high such as COLA1 are rare, and this might indicate that these are a product of short-lived star-formation bursts which many galaxies undergo during the EoR (e.g., Endsley et al., 2023). In that case, COLA1 would have likely needed the ionizing photon input from other sources to carve the bubble, specially for . Depending on the star-formation burstiness, the contribution from numerous non-detected faint sources () might be needed (see e.g., Witstok et al., 2023). In turn, if bursty star-forming episodes are common, they could account for a significant part of the ionizing photon budget during reionization. It was found in simulations that the ionizing photon escape fraction of individual galaxies can rapidly fluctuate in Myr scales (e.g., Rosdahl et al. 2022; see also Matthee et al. 2022). In such scenarios, a relatively small fraction of relatively bright galaxies are responsible of the bulk of the ionizing emissivity.
@@ 6.4 Is COLA1 powered by an AGN?
While the properties of COLA1 display particular similarities with some of the most extreme compact starbursts (e.g. Vanzella et al., 2022; Marques-Chaves et al., 2022; Ferrara, 2023; Topping et al., 2024), the extreme ionizing conditions and the compact size could also indicate that COLA1s ionizing radiation originates from an AGN instead. The narrowness of the Ly line suggests that COLA1 is not powered by an AGN (see Matthee et al., 2018). We also look for hints of an AGN in the H line measured in our grism data. First, we fit a Gaussian profile to the nebular [O III]λ5008 line. Next, we fit a double component Gaussian to the H line in order to capture possible broad and narrow line components (see Matthee et al., 2023b, for a similar procedure). The FWHM of one of the components is fixed to be the same as the fitted FWHM for [O III]λ5008, and the same value is used as lower limit for the other component. The best-fit for the broad component is FWHM km/s, with a contribution of % from the broad component to the total flux (S/N). The best fitting parameters lead to an AGN bolometric luminosity of , which would be able to explain only of the UV luminosity for a typical AGN spectrum (Shen et al., 2020). This result points towards a sub-dominant contribution to the UV luminosity.
The measured ratios [O III] and [O III]λ5008 / [O III] of COLA1 are typically seen in low- star-forming galaxies, whilst AGNs present higher [O III] for a given electron temperature (Übler et al., 2023). Hence, this is further evidence that an AGN is not the main driver of the ISM conditions in COLA1.
@@ 6.5 Future directions
Further in-depth analysis of COLA1 and its surrounding environment is needed to determine the properties of the system and its ionized bubble with precision. The particular shape of the rest-frame UV SED of COLA and its extreme UV slope are a challenge for the SED fitting models employed in this work (Sect. 3.3). Deeper, extended UV spectroscopy would be useful, in particular to confirm the particular SED of COLA1 and to test other indirect indicators of (LyC). These include coverage of [O II] (e.g., Mascia et al., 2023b; Choustikov et al., 2023), Mg II (e.g., Chisholm et al., 2020; Leclercq et al., 2024) or the rest-frame UV (e.g., C IV; Naidu et al., 2022a; Schaerer et al., 2022). It would be particularly intruiging to identify possible strong emission-lines such as NIV or CIV (e.g. Bunker et al., 2023; Calabro et al., 2024; Topping et al., 2024) that could boost and impact the rest-frame UV photometry.
On the other hand, the study of the other detected galaxies in COLA1’s over-density could yield valuable information about the size, morphology and optical depth of the ionized bubble. A natural future step is to follow-up these sources spectroscopically, aiming for the Ly line. If some of these galaxies are Ly emitters and they are located inside the bounds of the ionized bubble, their Ly line profiles and luminosities could be extremely valuable to determine the bubble properties. The enhanced Ly transmission of the ionized bubble would allow us to see the blue wings of the Ly lines and therefore estimate Ly and LyC escape fractions (e.g. Bosman et al., 2022).
@@ 7 Summary
COLA1 has a fairly high UV luminosity , a very steep UV slope and we infer negligible dust attenuation from the Balmer decrement. The observed SED of COLA1 is unusual and the models used, despite being fairly flexible, are only marginally consistent with the measured UV properties. The metallicity () and electron temperature ( K) can be directly estimated thanks to the H (S/N ) and [O III]λ4363 (S/N ) measurements; both appearing as typical values for galaxies at similar redshifts in the literature (Sects. 4.2, 4.3 and 3.3, Table 2).
@@
The observations presented in this work simultaneously probed the properties of the unusual COLA1 galaxy (morphology, UV slope and optical lines) and those of galaxies within its environment. We have shown evidence of a luminous galaxy that is actively ionizing its surroundings by showcasing the main characteristics that are common in LyC leakers, in particular the strong Ly emission with a narrow peak separation and a high SFR surface density. Simultaneously, we have shown that COLA1 resides in an over-density of galaxies and that these galaxies are located within the estimated size of the ionized bubble. However, the over-density is very common for normal galaxies that have similar UV or [O III] luminosities as COLA1. This suggests that the detection of the blue peak of the Ly line in COLA1 is not due to an unusually large over-density of galaxies that could have led to an unusually large ionised bubble that facilitated the observability of these blue Ly photons. An implication is that care should be taken when interpreting over-densities around galaxies with detected Ly lines in connection with reionization models. Now, the key next step in determining the impact of similarly luminous galaxies for cosmic reionization is to evaluate how common galaxies with Ly profiles like COLA1 are and what is the range in their over-densities.

### v2

@@ 5 The environment of COLA1: a normal over-density for its UV and [O III] luminosity
During the EoR, galaxy over-densities are associated with an increased Ly transmission (e.g., Endsley & Stark, 2022; Leonova et al., 2022; Kashino et al., 2023; Witstok et al., 2024), as these over-densities are more likely to form ionized bubbles, according to simulations (Ocvirk et al., 2020; Hutter et al., 2021; Qin et al., 2022; Lu et al., 2023). As sight-lines that allow for the detection of blue Ly photons at are expected to be extremely rare (e.g., Gronke et al., 2021), one could therefore expect that galaxies that show a double-peaked Ly line are surrounded by an unusually high over-density.
To address exactly this question, we analyzed the over-density around COLA1 and compared it with results from mock observations of a simulation, with the methodology we described in Sect. 3.5. We identify five [O III] emitters in our sample that are close to COLA1 in redshift space (; three of which have erg s-1). These galaxies are all considerably fainter than COLA1, spanning UV luminosities M to M. Notably, as is shown in Fig. 11 the sky position of all those objects fall within COLA1’s inferred ionized bubble radius (Matthee et al., 2018; Mason & Gronke, 2020), assuming a spherically symmetrical bubble, meaning that they could potentially be located within the spatial extent of the bubble. By comparing the number of identified sources with the expectation from pure random mocks (Sect. 3.5), we find that COLA1 is located in a mild over-density (). The typical halo mass for galaxies in our mocks that are surrounded by a similar number of galaxies as COLA1 is .
@@
This analysis shows that, while we do identify an over-density around COLA1, its amplitude is typical for any galaxy that is similarly UV bright at . Thus, there is no strong indication that the detection of the double-peaked Ly emission is due to an exceptionally large over-density of galaxies. Also, we remark that Fig. 12 shows that the possible number of neighbors that one could detected around a similarly UV bright galaxy with a similar emission-line survey on a field of arcmin2 has a wide distribution. This displays the wide variety of environments that could host similarly UV bright (; see e.g., Bouwens et al., 2015) galaxies, and therefore stressing the importance of taking cosmic variance of the large scale structure into account when interpreting Ly observations of galaxies in the context of reionization.
@@ 6 Discussion
In this section, we discuss the results presented in Sections 4 and 5. We comment the measured properties of COLA1 in comparison with other SFGs and LyC leakers in the literature (Sect. 6.1). Our results confirm the systemic redshift of COLA1, which means it must reside inside an ionized bubble (Sect. 6.2). We discuss the role of COLA1 and the galaxies in its proximity in ionizing their environment (Sect. 6.3) and we address the possible contribution of an AGN to COLA1’s luminosity (Sect. 6.4). Finally, we comment on some directions that can be followed in future works to further investigate the nature of COLA1 (Sect. 6.5).
@@ 6.1 COLA1 as an ionizing photon leaker
COLA1 presents an extremely blue UV slope (see Table 2), far bluer than the typical values of the C1F+EIGER sample and the best fit in Bouwens et al. (2014) for Lyman-break galaxies at (see Fig. 8). Topping et al. (2024a) presented a sample of 44 galaxies with steep UV slopes (), and found that the SEDs of such galaxies could be explained with density-bounded models with high escape fractions () (see also Topping et al., 2022). Further, Kim et al. (2023) found a relation between very steep and LyC leaking regions of a lensed galaxy. Indeed, our estimation based on Chisholm et al. (2022) using suggests a very high (see Table 3).
The rest-frame [O III]+H EW of COLA1 is comparable to the typical values for galaxies with similar magnitudes in UV-selected (e.g., Endsley et al., 2023a) and [O III] selected samples (e.g., Matthee et al., 2023). The intrinsic [O III] and H luminosities are remarkably high, and . The high [O III] luminosity can be related to an elevated star formation (Villa-Vélez et al., 2021), as well as the high ratio of (Dickey et al., 2016).
COLA1 presents significantly higher than the average C1F+EIGER values (see Fig. 9), consistent with the measurements for high- SFGs in the literature (e.g., Bouwens et al., 2016; Nakajima et al., 2018; Saxena et al., 2024; Simmonds et al., 2023, 2024) and also known low- LyC leakers (e.g., Izotov et al., 2016). Izotov et al. (2021) showed that low- compact SFGs are likely analogs of high- SFGs. More recently, compact SFGs have been found at with similar sizes to COLA1, presenting signs of LyC leakage (Mascia et al., 2024, 2023a).
As stated in Sect. 4.5, the direct measurement of H allows for an estimation of the Ly escape fraction, yielding a significantly high value of (Ly) . This Ly escape fraction is high in comparison with the typical values found in high redshift surveys (e.g., (Ly) %, for Ly selected sample at , Roy et al. 2023; (Ly) %, , mean , Tang et al. 2023), although other objects with similarly high Ly escape fractions have recently been found for fainter systems ((Ly) %, , –, Chen et al. 2024; (Ly) %, , Saxena et al. 2023). Moreover, as already discussed in Sect. 4.5, COLA1 shows a relatively high , also suggestive of high LyC escape fraction.
Almost every observed LyC leaker in the literature presents higher than the average at similar redshifts (e.g., Naidu et al., 2020). This is also the case of COLA1. Comparably, the galaxies Ion2 and Ion3 show fairly higher than the average at their respective redshifts, also presenting complex, multi-peaked Ly profiles (Vanzella et al., 2016, 2018).
@@ 6.2 Size of the ionized bubble
The size of the implied ionized region in the line of sight to COLA1 can be derived by the cutoff in the blue wing of the Ly line at km s-1 (see Fig. 4). This cutoff is particularly abrupt compared to the red peak, and defines the the line-of-sight extent in which the IGM is optically thin to Ly photons (so-called proximity zone; Mason & Gronke, 2020). The blue-peak cutoff seen in the Ly profile of COLA1 yields a proximity zone of pMpc (Matthee et al., 2018). Mason & Gronke (2020) argue that the total size of the ionized region extends beyond the proximity zone as there may be significant damping wing absorption. They find that a bubble with a radius of at least pMpc is needed in order to explain the observed velocity offset of COLA1’s blue Ly peak, assuming a residual neutral fraction of and defining the minimum observable velocity offset where the IGM transmission of Ly photons drops to 10%.
As is discussed in Mason & Gronke (2020), for a central source to be able to ionize a bubble with the required size pMpc, it needs a steep UV slope (), large escape fraction of ionizing photons (%) and an under-dense line-of-sight gas. Our results show that COLA1 aligns with the first two conditions, but we do not find a strong indication for a specifically under-dense line-of-sight (see Sect. 5). However, the estimated UV slope of COLA1 is steeper than the values considered in Mason & Gronke (2020), and its extreme value could explain this discrepancy as it implies a higher ionizing photon production efficiency. A more detailed analysis of the ionization conditions of the bubble is needed in order to address this topic, for example using Ly line measurements from multiple neighboring galaxies and by detailed comparisons to radiative hydrodynamical simulations (e.g., Gronke et al., 2021), which is beyond the scope of this work. In Sect. 6.3 below we discuss the role of COLA1 in ionizing such bubble.
@@ 6.3 COLA1 ionizing its surroundings
We investigate the contribution from COLA1 and the surrounding galaxies in the C1F group to the ionization of a pMpc bubble needed to explain COLA1’s Ly profile (see Sect. 6.2). The capability of galaxies of ionizing their medium directly depends on their LyC escape fraction. We estimate (LyC) for our sample with F200W detection (4 out of those 5 objects are detected in this band) using the relation with (UV) in Naidu et al. (2020), as is described in Sect. 4.5, for which we get (LyC) and %. For these estimates we have assumed the average dust correction introduced in Sect. 6.1.
@@
Equation 6 in Mason & Gronke (2020) predicts the radius of an ionized bubble powered by a central source as a function of an age parameter for a simple spherical geometry around the ionizing source. Using this equation, and following Witstok et al. (2024), we compute the radius of an ionized bubble powered by COLA1 for different values of the LyC escape fraction (Fig. 13). We have assumed an average hydrogen density cm-3 (Mason & Gronke, 2020) and an ionizing photon production of s-1 (Schaerer et al., 2016). For accounting the contribution from the other galaxies, in Fig. 13 we have conservatively assumed a high fiducial value of (LyC) 999For the computation of for the galaxies other than COLA1 we adopted ( Hz erg-1; Saxena et al., 2024), since a detection of H is not available for some of the objects.. According to this estimate, COLA1 alone could ionize a 0.7 pMpc bubble in a reasonable span of time ( Myr; see Witstok et al., 2024; Whitler et al., 2024) for escape fractions as low as , assuming it maintained the observed star-formation activity. The other four detected galaxies at make only a minor contribution in ionizing the bubble when combined with COLA1, specially for high COLA1 escape fractions (dotted lines in Fig. 13). For instance, for COLA1 LyC escape fraction of % (30%, 20%), the time required to ionize the bubble is Myr (58, Myr), with COLA1 contributing to (87%, ) of the total ionizing photon budget of the system detected in C1F.
@@
Nevertheless, systems with extremely high such as COLA1 are rare, and this might indicate that these are a product of short-lived star-formation bursts which many galaxies undergo during the EoR (e.g., Endsley et al., 2023b). In that case, COLA1 would have likely needed the ionizing photon input from other sources to carve the bubble, specially for . Depending on the star-formation burstiness, the contribution from numerous undetected faint sources () might be needed (see e.g., Witstok et al., 2024). In turn, if bursty star-forming episodes are common, they could account for a significant part of the ionizing photon budget during reionization. It was found in simulations that the ionizing photon escape fraction of individual galaxies can rapidly fluctuate in Myr scales (e.g., Rosdahl et al. 2022; see also Matthee et al. 2022). In such scenarios, a relatively small fraction of relatively bright galaxies are responsible of the bulk of the ionizing emissivity.
@@ 6.4 Is COLA1 powered by an AGN?
While the properties of COLA1 display particular similarities with some of the most extreme compact starbursts (e.g., Vanzella et al., 2022; Marques-Chaves et al., 2022; Ferrara, 2024; Topping et al., 2024b), the extreme ionizing conditions and the compact size could also indicate that COLA1s ionizing radiation originates from an AGN instead. The narrowness of the Ly line suggests that COLA1 is not powered by an AGN (see Matthee et al., 2018). We also look for hints of an AGN in the H line measured in our grism data. First, we fit a Gaussian profile to the nebular [O III]λ5008 line. Next, we fit a double component Gaussian to the H line in order to capture possible broad and narrow line components (see Matthee et al., 2024, for a similar procedure). The FWHM of one of the components is fixed to be the same as the fitted FWHM for [O III]λ5008, and the same value is used as lower limit for the other component. The best-fit for the broad component is FWHM km/s, with a contribution of % from the broad component to the total flux (S/N). The best fitting parameters lead to an AGN bolometric luminosity of , which would be able to explain only of the UV luminosity for a typical AGN spectrum (Shen et al., 2020). This result points towards a subdominant contribution to the UV luminosity.
The measured ratios [O III] and [O III]λ5008 / [O III] of COLA1 are typically seen in low- star-forming galaxies, whilst AGNs present higher [O III] for a given electron temperature (Übler et al., 2024). Hence, this is further evidence that an AGN is not the main driver of the ISM conditions in COLA1.
@@ 6.5 Future directions
Further in-depth analysis of COLA1 and its surrounding environment is needed to determine the properties of the system and its ionized bubble with precision. The particular shape of the rest-frame UV SED of COLA and its extreme UV slope are a challenge for the SED fitting models employed in this work (Sect. 3.4). Deeper, extended UV spectroscopy would be useful, in particular to confirm the particular SED of COLA1 and to test other indirect indicators of (LyC). These include coverage of [O II] (e.g., Mascia et al., 2023a; Choustikov et al., 2024), Mg II (e.g., Chisholm et al., 2020; Leclercq et al., 2024) or the rest-frame UV (e.g., C IV; Naidu et al., 2022a; Schaerer et al., 2022). It would be particularly intruiging to identify possible strong emission-lines such as NIV or CIV (e.g., Bunker et al., 2023; Calabro et al., 2024; Topping et al., 2024b) that could boost and impact the rest-frame UV photometry.
On the other hand, the study of the other detected galaxies in COLA1’s over-density could yield valuable information about the size, morphology and optical depth of the ionized bubble. A natural future step is to follow-up these sources spectroscopically, aiming for the Ly line. If some of these galaxies are Ly emitters and they are located inside the bounds of the ionized bubble, their Ly line profiles and luminosities could be extremely valuable to determine the bubble properties. The enhanced Ly transmission of the ionized bubble would allow us to see the blue wings of the Ly lines and therefore estimate Ly and LyC escape fractions (e.g., Bosman et al., 2022).
@@ 7 Summary
COLA1 has a fairly high UV luminosity , a very steep UV slope and we infer negligible dust attenuation from the Balmer decrement. The observed SED of COLA1 is unusual and the models used, despite being fairly flexible, are only marginally consistent with the measured UV properties. The metallicity () and electron temperature ( K) can be directly estimated thanks to the H (S/N ) and [O III]λ4363 (S/N ) measurements; both appearing as typical values for galaxies at similar redshifts in the literature (Sects. 4.2, 4.3 and 3.4, Table 2).
@@
The observations presented in this work simultaneously probed the properties of the unusual COLA1 galaxy (morphology, UV slope and optical lines) and those of galaxies within its environment. We have shown evidence of a luminous galaxy that is actively ionizing its surroundings by showcasing the main characteristics that are common in LyC leakers, in particular the strong Ly emission with a narrow peak separation and a high SFR surface density. Simultaneously, we have shown that COLA1 resides in an over-density of galaxies and that these galaxies are located within the estimated size of the ionized bubble. However, the over-density is very common for normal galaxies that have similar UV or [O III] luminosities as COLA1. This suggests that the detection of the blue peak of the Ly line in COLA1 is not due to an unusually large over-density of galaxies that could have led to an unusually large ionized bubble that facilitated the observability of these blue Ly photons. An implication is that care should be taken when interpreting over-densities around galaxies with detected Ly lines in connection with reionization models. Now, the key next step in determining the impact of similarly luminous galaxies for cosmic reionization is to evaluate how common galaxies with Ly profiles like COLA1 are and what is the range in their over-densities.